OpenStudy (hitaro9):

Convert to Cartesian Coordinates? r= sin(2theta)

OpenStudy (anonymous):

do you know the answer? mine is coming out goofy

OpenStudy (hitaro9):

I know that a squareroot is involved. Webassign says "Right the answer so any square root has a positive coefficient"

OpenStudy (hitaro9):

Write*

OpenStudy (anonymous):

i put \(\sin(2\theta)=2\cos(\theta)\sin(\theta)\) and got \[r=2\sin(\theta)\cos(\theta)\] then replaced \(\cos(\theta)\) by \(\frac{x}{r}\) and \(\sin(\theta)\) by \(\frac{y}{r}\) and got \[r=2\frac{xy}{r^2}\] then \[r^3=2xy\] or \[(x^2+y^2)\sqrt{x^2+y^2}=2xy\] but that looks goofy to me

OpenStudy (hitaro9):

I'm seeing the logic, but yeah. Most of the other answers have worked out very well and looked a lot simpler.

OpenStudy (anonymous):

think i missed at \(\pm\)

OpenStudy (hitaro9):

I'm gunna go ahead and enter what you put in to webassign. If it works out, then I guess it was just a complicated answer.

OpenStudy (anonymous):

gotta love webassign good luck!

OpenStudy (hitaro9):

Yep. It worked out. Thanks a ton dude.